# LCEN: A Novel Feature Selection Algorithm for Nonlinear, Interpretable Machine Learning Models

### v1

@@ 3.1 Artificial data highlight LCEN’s robustness to noise, multicollinearity, and hyperparameter variance
On this same dataset and using the same four types of models, LCEN attained median errors that are typically over 60% smaller than for ALVEN (Figure 2). Similarly to the models generated using ALVEN, the model with a degree between 1 and 10 “cv” had the lowest performance and the model with degree equal to 2 or 4 “cv limited order” had a performance between the degree 4 “unbiased model” and the degree 2 “biased model”. However, the degree 4 “unbiased model” was always the best model, no matter the noise level used. We attribute this considerable reduction in median test MSEs and the superiority of the degree 4 “unbiased model” created by LCEN to the improved feature selection algorithm, which is able to better resist variance due to noise and a large number of hyperparameters. This is corroborated by how the model with a degree between 1 and 10 “cv” tended to select degree = 4 at lower noise levels and degree = 2 at higher noise levels (Figure A1), showing how LCEN can automatically follow the hypothesis set by Sun & Braatz (2021) regarding the bias-variance tradeoff.
@@
Figure 2: Test set median MSE for the “4th-degree, univariate polynomial” dataset. ALVEN results (left, reproduced from Sun & Braatz (2021) with permission) show that the error is monotonically increasing with noise and that the degree 4 “unbiased model” is the best at low noise levels, but is displaced by the degree 2 “biased model” at higher noise levels. On the other hand, LCEN results (right) show that the median errors converge at higher noises. Furthermore, the LCEN median errors are typically over 60% smaller than the ALVEN median errors, and the degree 4 “unbiased model” is always the best model no matter the noise. The “noise level” and “Noise variance $\sigma^{2}$” terms are equivalent in this figure.
3.2 Empirical data from processes with known physical laws further highlight LCEN’s feature selection potential
The applicability of an algorithm to real-world problems is judged only by its performance on real data, as data sparsity or noise may affect the algorithm’s capabilities. These tests are summarized in Table A3. The first test uses the “CARMENES star data” dataset from Schweitzer et al. (2019). This dataset contains information on temperature ($T$), radius ($R$), and luminosity ($L$) of 293 white dwarf stars. These features are linked together by the Stefan-Boltzmann equation, $L=4\pi R^{2}\sigma T^{4}$, where $\sigma$ is a constant. Normalizing this equation to values from another star (typically, the Sun), conveniently sets the constant terms to 1. This normalization is applied to the “CARMENES star data” dataset. LCEN with degrees from 1 to 10 was applied to this normalized dataset. Despite the very large number of potential features (due to the high degree values used), LCEN correctly selected only the $R^{2}T^{4}$ feature. The coefficient assigned to $R^{2}T^{4}$ is 0.9826, which is well within the 2–3% error on these data (as reported by Schweitzer et al. (2019)). LCEN retained high performance for this real data in a high-hyperparameter variance scenario.
A potential limitation in real datasets is data scarcity. To evaluate the LCEN algorithm in a low-data scenario, the “Kepler’s 3rd Law” datasets are created. The first version uses the original data obtained by Kepler, first published in 1619 and republished in Kepler et al. (1997). From only 6 (slightly inaccurate) measurements, Kepler was able to derive the eponymous Kepler’s 3rd law, which states that the period $T$ of a celestial body is related to the semi-major axis of its orbit $a$ by $T=ka^{3/2}$. The constant $k$ depends on the masses of the central and orbiting bodies; however, as the mass of the central body is typically much larger, the mass of the orbiting body is ignored. In this and Kepler’s works, $T$ is measured in Earth days, so the constant $k$ is $\sim$365.25 when using modern data and $\sim$365.15 when using Kepler’s original data. Despite the low number of data points, LCEN correctly selected only the $a^{3/2}$ feature. Moreover, the coefficient assigned to that feature was 366.82, an error of only 0.46% relative to Kepler’s $k=365.15$.
LCEN is then evaluated using a modern version of the same dataset, which contains 8 points (as Uranus and Neptune were discovered after Kepler’s observations) whose data were measured with greater accuracy. On this modern “Kepler’s 3rd Law” dataset, LCEN again selects only the $a^{3/2}$ feature. The coefficient assigned to the $a^{3/2}$ feature is 365.00, an error of only 0.07% relative to the modern value $k=365.25$. LCEN did perfect feature selection in these data-scarce scenarios, with parameter estimates minimally affected by experimental noise.
Lastly, LCEN is tested on the “Newton’s Law of Cooling” dataset to determine its ability to work with scarce, time-series data approximated from a differential equation. This dataset was published by Newton (1701) and republished in English by Simms (2004). Newton used 14 measurements to determine the cooling rate of a “pretty thick piece of red-hot iron” (Simms, 2004). While Newton made some improper assumptions in that work (including ignoring thermal radiation emissions, which are calculated through the Stefan-Boltzmann equation), his measurements are accurate for $T\leq 300^{\circ}$ (Simms, 2004). In its non-dimensionalized form, Newton’s Law of Cooling relates the rate of change of temperature of an object to its temperature $T(t)$ and the temperature of its surroundings $T_{\text{env}}$, such that $\frac{dT}{dt}=rT_{\text{env}}-rT(t)$. The constant $r$ is the heat transfer coefficient which depends on the object’s mass, surface area $A_{s}$, specific heat capacity $c_{p}$, and fluid heat transfer coefficient $h$, such that $r=\frac{hA_{s}}{mc_{p}}$. Applying a first-order finite difference approximation and rearranging replaces Newton’s Law of Cooling by $T(t)=rT_{\text{env}}\Delta t+T(t-1)(1-r\Delta t)$.
[…]
@@
3.3 LCEN surpasses many other models when making predictions on empirical data from processes with unknown physical laws
@@
The final experiments to validate LCEN’s performance involve comparisons to other algorithms on real datasets from processes with unknown physical laws. As there is no (computational) way to validate the feature selection by models trained on these datasets, the main focuses of this section are investigating prediction errors and sparsities of different model architectures.
@@
The first dataset analyzed is the “Diesel Freezing Point” dataset (Hutzler & Westbrook, 1999), which is comprised of 395 diesel spectra measured at 401 wavelengths and used to predict the freezing point of these diesels. LCEN; the dense model architectures of ordinary least squares (OLS), ridge regression (RR), and partial least squares (PLS); and the sparse model architectures LASSO and EN are compared in Table 3. Most models had test set RMSEs slightly smaller than ${}^{\circ}$, which is about 7.5% of the range of the test data, which contains diesels with freezing points between $-59.5^{\circ}$ and $6.6^{\circ}$. RR and EN had the lowest test RMSEs of ${}^{\circ}$, and LCEN followed right after with a test RMSE of $4.86^{\circ}$. However, the best hyperparameters for EN included an l1_ratio = 0.1, which makes its model very dense. Whereas RR selected all 401 features and EN selected 299, LCEN was able to achieve similar performance with only 37 features. Moreover, LCEN is also faster: whereas EN took 19.6 s to run, LCEN took only 6.82 s. Finally, the LCEN cutoff hyperparameter was increased from the value that minimizes the validation MSE to create sparser models. These models have much fewer variables, yet their test set RMSEs increase by less than ${}^{\circ}$. This illustrates how LCEN can select the most critical features to make models with high sparsity and predictive power, and how these criteria can be prioritized by the end-user.
@@
Architecture Test RMSE (${}^{\circ}$) Features Runtime (s) OLS 11.75 401 0.09 PLS 5.21 401 4.44 RR 4.83 401 4.87 EN 4.83 299 19.6 LASSO 4.90 39 2.06 LCEN 4.86 37 6.82 4.95 20 6.43 5.16 10 6.02 5.47 5 5.82 5.80 3 5.85
@@
The next dataset used is the “Abalone” dataset (Nash et al., 1995). Abalone (Haliotis sp.) are sea snails whose age can be determined by cutting their shells, staining them, and counting the stained shell rings under a microscope. This process is laborious and error-prone. An alternative is to estimate the number of rings based on readily available physical characteristics, such as weight and size. As before, LCEN was compared with other dense and sparse machine learning models, and LCEN models with increased sparsity were also generated (Table 4). In this problem, OLS, RR, LASSO, and EN all converged to the OLS solution (that is, no regularization), selecting all 9 linear features and having an RMSE of 2.1 rings. On the other hand, LCEN automatically detected that 2nd degree features would be relevant. The model with the LCEN algorithm selected 11 features (both 1st- and 2nd-degree features) and had an RMSE of 2.0 rings. By increasing the sparsity, an LCEN model with only 3 features (one of which was a 2nd-degree feature) had an RMSE of 2.2 rings. This experiment further illustrates LCEN’s selecting the best features even when hyperparameter variance is present, and very sparse LCEN models retaining significant performance.
Table 4: Results of different model architectures for the “Abalone” dataset.
Architecture Test RMSE (rings) Features OLS=RR=LASSO=EN 2.1 9 LCEN 2.0 11 2.2 3
@@
LCEN is then tested on the “Concrete Compressive Strength” dataset (Yeh, 1998), which contains the composition and age of 1,030 different types of concrete and their compressive strengths. The relationship between these properties is nonlinear, and previous modeling attempts include algebraic expressions and artificial neural networks (Yeh, 1998, 2006). These ANNs were superior to the algebraic models, whereas the algebraic models provide interpretability on how the properties of the concrete affect its compressive strength. LCEN is also considerably better than the previously published algebraic models (Table 5), and its performance is competitive with that of previously published ANNs without sacrificing interpretability. Furthermore, note that no type of validation is mentioned in Yeh (1998), so the test and validation sets may be the same, making the ANN figures overoptimistic. LCEN has a validation RMSE of 4.78 MPa and an $R^{2}$ of 0.920 on this dataset, which are effectively equal to those from the ANN of Yeh (1998).
@@
Table 5: Results of different model architectures for the “Concrete Compressive Strength” dataset. No type of validation is mentioned in Yeh (1998), so the test and validation sets may be the same, making the ANN values overoptimistic.
[…]
@@ 4 Discussion
This work introduces the LASSO-Clip-EN (LCEN) algorithm for the creation of nonlinear, interpretable machine learning models (Algorithm 1). The algorithm is first validated using artificial data (Section 3.1), which provide an initial assessment of the algorithm’s performance under different conditions, which can be controlled independently. LCEN displayed high performance for data that are noisy (“Linear, 5-variable polynomial” and “4th-degree, univariate polynomial” datasets; Table 1 and Fig. 2), multicollinear (“Multicollinear data” dataset; Fig. 1), or generated from a process with a complex nonlinear function (“Relativistic energy” and “4th-degree, univariate polynomial” datasets; Table 2 and Fig. 2), which also leads to high hyperparameter and feature variance. For the “4th-degree, univariate polynomial” dataset, LCEN automatically adjusted its predictions to compensate for the bias-variance tradeoff (Figs. 2 and A1). During the experiments with the “Relativistic energy” dataset, the range of the data affected the prediction performance – data with higher ranges led to more accurate predictions even at high noises. We hypothesize that this phenomenon occurs because it is difficult to distinguish between multiple powers of a variable (or, in a more generalized manner, how Taylor approximations are successful) when the data cover low ranges and are noisy.
@@
LCEN was then tested with data from processes with known physical laws (Section 3.2). These data have real properties (such as noise) and may have limitations common to many processes (such as low amounts of data). Nevertheless, LCEN successfully selected the correct features for all datasets used in this work, effectively rediscovering physical laws solely from data (Table A3). The first dataset tested was the “CARMENES star data”, which contains hundreds of points described by Stefan-Boltzmann law. This law uses a 6th-order interaction, but since this would not be known a priori, LCEN was tested using all degree hyperparameters from 1 to 10. Despite the high variance of the process and the many potential features that could be selected, LCEN selected only the correct $R^{2}T^{4}$ feature with a coefficient relative error of only 1.74%, which is well within the 2–3% data error. Next, LCEN was evaluated for the “Kepler’s 3rd Law” datasets, which contain small amounts of data (only 6 or 8 points) yet were enough for Kepler to derive the eponymous law in 1619. LCEN can automatically replicate this discovery despite the slightly inaccurate data from Kepler’s measurements, selecting only the correct $a^{3/2}$ feature. The coefficient relative errors were 0.46% on Kepler’s original data and 0.07% on modern data, extremely small values that highlight the algorithm’s accuracy. LCEN algorithm was also demonstrated for (scarce) time-series data, to model the “Newton’s Law of Cooling” dataset. LCEN again discovered the correct features (although a very small intercept was ignored), and the predicted coefficients were realistic according to estimations.
@@
Lastly, experiments using empirical data from processes with unknown laws further validated LCEN’s feature selection and predictive performance. LCEN was compared to multiple different model architectures, including LASSO, EN, and ANNs. On the “Diesel Freezing Point” dataset (Table 3), LCEN had a test RMSE within 1% of the RMSEs for RR and EN, which tied for the lowest RMSE. However, LCEN required 10.8 and 8.1 times fewer features than RR and EN (respectively) to reach the same accuracy, and LCEN was 2.9-fold faster than EN. LCEN also generated a more accurate and sparser model than LASSO, and LCEN models with increased sparsities displayed only small increases in the prediction error. In the most extreme case, an LCEN model that used only 3/401 features had a test RMSE less than ${}^{\circ}$ larger than the best models. LCEN provided a combination of accuracy, sparsity/interpretability, and speed that was unmatched by the traditional dense and sparse linear machine learning methods. On the “Abalone” dataset (Table 4), LCEN automatically recognized the importance of 2nd-degree features, which led to its producing the model with the lowest RMSE. OLS, RR, LASSO, and EN converged to the same dense linear model. Similarly to what happened with the “Diesel Freezing Point” dataset, an LCEN model using only 3 features, one of which was 2nd-degree, displayed only a 2.2% increase in RMSE relative to the dense models. As before, LCEN built very sparse yet very accurate models. Finally, LCEN was compared to an algebraic expression, a machine learning model, and an ANN on the “Concrete Compressive Strength” dataset (Table 5). LCEN displayed a test RMSE at least 22% lower than that of the other interpretable models and produces a model that is simpler and interpretable while having an RMSE competitive with that of the ANN. Yeh (1998), the paper that originated the ANN, does not mention any type of validation, which could indicate that the ANN results are overoptimistic. The validation performance of LCEN was equal to the ANN test performance reported in Yeh (1998), indicating LCEN is competitive even with ANNs on this problem if the results reported by Yeh (1998) are actually validation results.
@@
Overall, these experiments have demonstrated the applicability of LCEN to a multitude of scientific problems. The models trained by LCEN were robust to defects in the real data, including noise, multicollinearity, or sample scarcity. LCEN models were as accurate as or more accurate than many alternative architectures, yet were also much sparser. This combination of accuracy and interpretability is essential for the deployment of machine-learning models in performance-critical scenarios, from aviation to medicine. LCEN is free, open-source, and easy to use, allowing even non-specialists in machine learning to benefit from the algorithm and use it in technical work. Moreover, the additional interpretability can assist in data or model refinement efforts and can make the models robust to changes in data or adversarial input. Finally, this mathematical, feature-based algorithm can also be used to automate efforts in certain hybrid model architectures (such as physics-constrained or physics-guided ML).
Impact Statement
The main societal impact of this paper is the creation of interpretable and accurate machine learning models. These may increase the robustness, trust, and reliability of ML models applied in many scenarios, including performance-critical scenarios such as aviation and medicine. A discussion of the benefits of interpretability is beyond the scope of this section; many other works have been dedicated to this, such as Rudin (2019).

### v2

@@ 3.1 Artificial data highlight LCEN’s robustness to noise, multicollinearity, and hyperparameter variance
On this same dataset and using the same four types of models, LCEN attained median errors that are typically over 60% smaller than for ALVEN (Figure 1). Similarly to the models generated using ALVEN, the model with a degree between 1 and 10 “cv” had the lowest performance and the model with degree equal to 2 or 4 “cv limited order” had a performance between the degree 4 “unbiased model” and the degree 2 “biased model”. However, the degree 4 “unbiased model” was always the best model, no matter the noise level used. We attribute this considerable reduction in median test MSEs and the superiority of the degree 4 “unbiased model” created by LCEN to the improved feature selection algorithm, which is able to better resist variance due to noise and a large number of hyperparameters. This is corroborated by how the model with a degree between 1 and 10 “cv” tended to select degree = 4 at lower noise levels and degree = 2 at higher noise levels (Figure A3), showing how LCEN can automatically follow the bias-variance tradeoff hypothesis.
@@
Figure 1: Test set median MSE for the “4th-degree, univariate polynomial” dataset. ALVEN results (left, reproduced from Ref. [35] with permission) show that the error is monotonically increasing with noise and that the degree 4 “unbiased model” is the best at low noise levels, but is displaced by the degree 2 “biased model” at higher noise levels. On the other hand, LCEN results (right) show that the median errors converge at higher noises. Furthermore, the LCEN median errors are typically over 60% smaller than the ALVEN median errors, and the degree 4 “unbiased model” is always the best model no matter the noise. The “noise level” and “Noise variance $\sigma^{2}$” terms are equivalent in this figure. Interquartile ranges for the LCEN model’s test MSEs are available in Fig. A2.
@@
3.2 LCEN surpasses many other architectures when making predictions on empirical data
The applicability of an algorithm to real-world problems is judged only by its performance on real data, as data sparsity or real noise may affect the algorithm’s capabilities. Tests done on empirical data generated by processes with known physical laws show that LCEN still displays exceptional feature selection capabilities, consistently selecting only the right features even when high hyperparameter variance is present (Section A6.1 and Table 3).
Table 3: Summary of LCEN results for the empirical datasets from processes with known physical laws.
Dataset Only Correct Features Relative Error of the Predicted Coefficient CARMENES star data Yes 1.74% Kepler’s 3rd Law (1619) Yes 0.46% Kepler’s 3rd Law (Modern) Yes 0.07%
@@ 3.2 LCEN surpasses many other architectures when making predictions on empirical data
The final experiments to validate LCEN’s performance involve comparisons to other algorithms on real datasets from processes with unknown physical laws. As there is no (computational) way to validate the feature selection by models trained on these datasets, the main focuses of this section are investigating prediction errors and sparsities of different model architectures. The linear methods ordinary least squares (OLS), ridge regression (RR) [36], partial least squares (PLS) [37, 38], LASSO, and elastic net (EN); the nonlinear ensemble methods random forest (RF) [39], gradient-boosted decision trees (GBDT) [40], adaptive boosting (AdaB) [41]; and the nonlinear methods support vector machine with radial-basis functions (SVM) [42], multilayer perceptron (MLP), MLP with group LASSO (MLP-GL1) [23], and LassoNet [25] were compared with LCEN. To clarify our design choices and the relevance of each individual part of the LCEN architecture, ablation tests are performed with many of the datasets tested here in Section A4 of the Appendix.
@@
The first dataset analyzed is the “Diesel Freezing Point” dataset [43], which is comprised of 395 diesel spectra measured at 401 wavelengths and used to predict the freezing point of these diesels. The dense, nonlinear architectures SVM and MLP had the best prediction performance, with test RMSEs equal to 4.39 and 4.61 C respectively (Table 4). They were followed by RR, EN, LCEN, LassoNet, LASSO, and MLP-GL1, which had test RMSEs between 4.83 and 4.92 C. This set of architectures include both linear and nonlinear methods, and both dense and sparse methods. Other model architectures performed worse, with test RMSEs > 5.0C. For comparison, 5.0C is about 7.5% of the range of the test data, which contains diesels with freezing points between $-59.5^{\circ}$ and $6.6^{\circ}$. The sparsest architecture was LCEN, which selected only 36/401 features (9.0%) yet had a prediction error only 10.0% higher than that of the best dense method. LCEN and SVM were the only nonlinear methods that had a runtime faster than 10 seconds, a speed typically reserved for linear methods. EN and AdaB had runtimes $\approx$ 20 seconds, and all other models had runtimes on the order of hundreds or thousands of seconds. LCEN is the only architecture that combines a low test RMSE, interpretability, and a fast runtime. LASSO, the only other model architecture that also has these properties, has a worse RMSE and sparsity. Finally, the LCEN cutoff hyperparameter was increased from the value that minimizes the validation MSE to create sparser models. These models have much fewer variables, yet their test set RMSEs typically increase by only small values. This illustrates how LCEN can select the most critical features to make models with high sparsity and predictive power, and how these criteria can be prioritized by the end-user.
@@
Architecture Test RMSE (C) Features Runtime (s) OLS 11.75 401 0.09 PLS 5.21 401 4.44 RR 4.83 401 4.87 EN 4.83 299 19.6 LASSO 4.90 39 2.06 RF 5.16 390 307 GBDT 5.40 354 2649 AdaB 5.60 300 22.0 SVM 4.39 401 5.33 MLP 4.61 401 121 MLP-GL1 4.92 80 569 LassoNet 4.83 401 474 LCEN 4.83 36 6.54 4.91 29 6.27 5.52 13 5.76 7.40 6 5.53
@@
The next dataset used is the “Abalone” dataset [44]. Abalone (Haliotis sp.) are sea snails whose age can be determined by cutting their shells, staining them, and counting the stained shell rings under a microscope. This process is laborious and error-prone. An alternative is to estimate the number of rings based on readily available physical characteristics, such as weight and size. As before, LCEN was compared with other dense and sparse machine learning models, and LCEN models with increased sparsity were also generated (Table A7). In this problem, OLS, PLS, RR, LASSO, and EN all converged to the OLS solution (that is, no regularization), selecting all 8 linear features and having an RMSE of 2.1 rings. On the other hand, LCEN automatically detected that 2nd degree features would be relevant. The model with the LCEN algorithm also selected all 8 features and had an RMSE of 2.0 rings. Nonlinear models had test RMSEs between 2.0 and 2.3 rings, but they all lack the interpretability of LCEN. By increasing the cutoff hyperparameter, sparser LCEN models may be generated. An LCEN model with only 3 features had an RMSE of 2.1 rings, and another with only 2 features had an RMSE of 2.2 rings. This experiment further illustrates LCEN’s robust feature selection, and how very sparse LCEN models retain significant performance.
@@
LCEN is then tested on the “Concrete Compressive Strength” dataset [34], which contains the composition and age of 1,030 different types of concrete and their compressive strengths. The relationship between these properties is nonlinear, and previous modeling attempts include algebraic expressions and artificial neural networks (specifically, MLPs) [34, 45]. These MLPs were superior to the algebraic models, whereas the algebraic models provide interpretability on how the properties of the concrete affect its compressive strength. LCEN is also considerably better than the previously published algebraic models (Table 5), and its performance is competitive with that of previously published ANNs without sacrificing interpretability. Furthermore, note that no type of validation is mentioned in [34], so the test and validation sets may be the same, making the MLP figures overoptimistic. LCEN has a validation RMSE of 4.66 MPa on this dataset, which is slightly better than that from the MLP of [34]. Some other machine-learning models surpass LCEN in terms of test RMSE, but LCEN has the lowest validation RMSE out of all architectures tested.
@@
Table 5: Results of different model architectures for the “Concrete Compressive Strength” dataset. All machine-learning models selected all 8 features. No type of validation is mentioned in Ref. [34], so the test and validation sets may be the same, making the ANN values overoptimistic.
Architecture Test RMSE (MPa) Algebraic expression [34] 7.79 MLP [34] 4.76 Linear+interactions model [45] 7.43 RF 5.10 GBDT 7.29 AdaB 6.95 SVM 5.94 MLP-GL1 5.47 LassoNet 5.53 LCEN 5.73
LCEN is also successful at predicting phenomena caused by human activity instead of physical laws. In the modified “Boston housing” dataset [46], LCEN attains a test RMSE that is only 6% higher than that of a dense MLP (Table A8). Once again, LCEN attains higher performance than many other architectures in this dataset while also being completely interpretable. Finally, the “GEFCom 2014” dataset was used to highlight the ability of LCEN to predict in a complex and dynamic task [47]. Two versions of the “GEFCom 2014” dataset have been published: one that contains only energy consumption levels and another that contains the same energy consumption data and also temperature data from multiple weather stations. This work uses the former. “GEFCom 2014” is part of an energy forecasting competition which was originally won by a LASSO-like model [47]. More recently, deep learning has been applied to this problem [48, 49], and deep learning models have achieved strong 24-hour predictive performance [49]. Despite the strong performance of multiple, complex ANN architectures, LCEN models obtain a 13.1% lower test RMSE on this forecasting task than the state-of-the-art Seq2Seq model from Ref. [49] (Table 6). Unlike the ANNs, LCEN requires only a CPU for training and forecasting, and provides interpretable coefficients. LCEN can also be used for longer forecasts without significant increases in the prediction error, further highlighting the robustness of the algorithm.
Table 6: Results of different model architectures for the “GEFCom 2014” dataset. The deep learning models (TCN to Seq2Seq) and their results come directly from Ref. [49].
Metric TCN RNN LSTM GRU Seq2Seq LCEN Hours Forecast 24 24 24 24 24 24 48 72 120 168 Mean Test RMSE (MW) 17.2 18.0 19.5 19.0 17.1 14.9 18.9 21.0 23.4 24.7 Mean Relative Error (%) 9.8 10.2 11.1 10.8 9.7 8.5 10.7 11.9 13.2 13.9
@@ 4 Discussion
This work introduces the LASSO-Clip-EN (LCEN) algorithm for the creation of nonlinear, interpretable machine learning models (Algorithm 1). LCEN is first validated using artificial data (Section 3.1), which provide an initial assessment of the algorithm’s performance under multiple, independently controllable conditions. LCEN displayed high performance for data that are noisy (“Linear, 5-variable polynomial” and “4th-degree, univariate polynomial” datasets; Table 1 and Fig. 1), multicollinear (“Multicollinear data” dataset; Fig. A1), or generated from a process with a complex nonlinear function (“Relativistic energy” and “4th-degree, univariate polynomial” datasets; Table 2 and Fig. 1), which also leads to high hyperparameter and feature variance. For the “4th-degree, univariate polynomial” dataset, LCEN automatically adjusted its predictions to compensate for the bias-variance tradeoff (Figs. 1, A2, and A3). During the experiments with the “Relativistic energy” dataset, the range of the data affected the prediction performance – data with higher ranges led to more accurate predictions even at high noises. We hypothesize that this phenomenon occurs because it is difficult to distinguish between multiple powers of a variable (or, in a more generalized manner, how Taylor approximations are successful) when the data cover low ranges and are noisy.
@@
LCEN was then tested with data from processes with known physical laws (Section A6.1). These data have real properties (such as noise) and may have limitations common to many processes (such as low amounts of data). Nevertheless, LCEN successfully selected the correct features for all datasets used in this work, effectively rediscovering physical laws solely from data (Table 3). The “CARMENES star data” dataset contains hundreds of points described by the Stefan-Boltzmann law. This law uses a 6th-order interaction, but since this would not be known a priori, LCEN was tested using all degree hyperparameters from 1 to 10. Despite the high variance of the process and the many potential features that could be selected, LCEN selected only the correct $R^{2}T^{4}$ feature with a coefficient relative error of only 1.74%, which is well within the 2–3% data error. Next, LCEN was evaluated for the “Kepler’s 3rd Law” datasets, which contain small amounts of data (only 6 or 8 points) yet were enough for Kepler to derive the eponymous law in 1619. LCEN can automatically replicate this discovery despite the slightly inaccurate data from Kepler’s measurements, selecting only the correct $a^{3/2}$ feature. The coefficient relative errors were 0.46% on Kepler’s original data and 0.07% on modern data, extremely small values that highlight the algorithm’s accuracy.
@@
Lastly, experiments using empirical data from processes with unknown laws further validated LCEN’s feature selection and predictive performance. LCEN was compared to multiple different model architectures, including linear models, ensemble methods, and sparse and dense MLPs. On the “Diesel Freezing Point” dataset (Table 4), LCEN had a test RMSE equal to the RMSEs for RR and EN. However, LCEN required 11.1- and 8.3-fold fewer features than RR and EN (respectively) to reach the same accuracy, and LCEN was 3.0-fold faster than EN. LCEN also generated a more accurate and sparser model than LASSO, and LCEN models with increased sparsities typically displayed only small increases in the prediction error. When compared to SVM, the architecture with the lowest test RMSE, LCEN displayed only a 10.0% higher error, despite being a completely interpretable architecture and using only 36/401 features. LCEN provided a combination of accuracy, sparsity, interpretability, and speed that was unmatched by the other machine-learning methods. On the “Abalone” dataset (Table A7), LCEN automatically recognized the importance of 2nd-degree features, which led to its producing the model with the lowest RMSE. All linear methods converged to the same dense model. Nonlinear methods had similar performance, reaching test RMSEs between 2.0 and 2.3 rings, showing LCEN can provide interpretability without sacrificing accuracy. Finally, an LCEN model using only 3 features displayed only a 2.3% increase in RMSE relative to the best LCEN model, and another LCEN model using only 2 features displayed only a 7.1% increase in RMSE relative to the same best LCEN model. As before, LCEN built very sparse yet very accurate models. LCEN was then compared to previously published algebraic expressions and an ANN, and many other machine-learning methods, on the “Concrete Compressive Strength” dataset (Table 5). LCEN displayed a test RMSE at least 23% lower than that of the algebraic models and produces a model that is simpler and interpretable while having an RMSE competitive with that of the MLP of Ref. [34]. In the paper that originated this MLP, no type of validation is mentioned, which could indicate that the MLP results are overoptimistic. The validation performance of LCEN was better than the ANN test performance reported in Ref. [34], indicating LCEN is competitive even with ANNs on this problem if the results reported by Ref. [34] are actually validation results. Some nonlinear machine-learning architectures achieved a lower test RMSE on this task, but LCEN has the lowest validation RMSE out of all architectures tested.
Continuing these experiments, LCEN was also tested against datasets generated by human activity, as opposed to physical phenomena. In a dataset used to predict the prices of houses (“Boston housing”), LCEN achieved strong performance, with a test RMSE only 6% higher than that of an MLP, which had the best test-set performance. This exceptional result further highlights the potential of LCEN, especially in data-deficient scenarios. In the “GEFCom 2014” dynamic dataset, LCEN was able to surpass complex ANN architectures, predicting test-set power loads with a 13.1% lower RMSE than that of the former state-of-the-art model (Table 6). Due to its high robustness, LCEN also could predict for longer time horizons without large increases in the prediction error.
@@
Overall, these experiments have demonstrated the applicability of LCEN to a multitude of scientific and nonscientific problems. LCEN models were robust to defects in the real data, including noise, multicollinearity, or sample scarcity. LCEN models were typically as accurate as or more accurate than many alternative architectures, yet were also much sparser. The LCEN architecture is also trivial to interpret and displays exactly how each input is contributing to the final output. This combination of accuracy and interpretability is essential for the deployment of machine-learning models in performance-critical scenarios, from aviation to medicine. LCEN is free, open-source, and easy to use, allowing even non-specialists in machine learning to benefit from and use it. Moreover, the additional interpretability can assist in data or model refinement efforts and can make the models robust to changes in data or adversarial input. The main limitations of LCEN are that it is not a universal function approximator, as it can model only the functions present in the expansion of dataset features, and that it sometimes is not as accurate as a dense deep learning method. If a GPU and enough time are available for model training, users in scenarios that focus on accuracy above anything else may prefer to use a deep learning architecture.
There are at least two clear future directions for this work. The first involves using the LCEN algorithm in classification tasks, as many important problems in science and engineering involve classification. A comprehensive and robust analysis of the performance of LCEN in classification tasks will follow this paper. A second future direction involves applying the LCEN algorithm to automatically generate physical equations for certain hybrid model architectures (such as physics-constrained or physics-guided ML). The potential and importance of these models has been analyzed in many works, such as in Refs. [50, 51].